OpenStudy (anonymous):

5/x^2 - 1 - 2/x = 2/x+1

hero (hero):

Remember to use parentheses 5/(x^2 - 1) - 2/x = 2/(x+1)

hero (hero):

Otherwise it can be interpreted differently

hero (hero):

Anyway.... \(\large\frac{5}{x^2-1} - \frac{2}{x} = \frac{2}{x+1}\)

hero (hero):

Make sure the denominators are factored: \(\large\frac{5}{(x+1)(x-1)} - \frac{2}{x} = \frac{2}{x+1}\)

hero (hero):

Multiply both sides by (x-1)

hero (hero):

\(\frac{5}{x+1} - \frac{2(x-1)}{x} = \frac{2(x-1)}{x+1}\)

hero (hero):

Add the second fraction to both sides; Subtract the third fraction from both sides: \(\frac{5}{x+1} - \frac{2x -2}{x+1} = \frac{2x-2}{x}\)

hero (hero):

Combine the fractions on the left hand side: \(\frac{5 - 2x + 2}{x+1} = \frac{2x-2}{x}\)

hero (hero):

Then simplify to get \(\frac{7 - 2x}{x+1} = \frac{2x-2}{x}\)

hero (hero):

Then cross multiply: x(7-2x) = 2(x-1)(x+1)

hero (hero):

\(7x-2x^2 = 2(x^2 - 1)\)

hero (hero):

\(7x - 2x^2 = 2x^2 - 2\)

hero (hero):

\(4x^2 - 7x - 2 = 0\)

hero (hero):

\(4x^2 -8x + x - 2 = 0 \\ 4x(x-2) + 1(x-2) = 0 \\ (x-2)(4x+1) = 0\) Continue Solving for x

hero (hero):

You should get: x = -1/4 x = 2

hero (hero):

Any questions?